The total accumulated costs C(t) and revenues R(t) (in thousands of dollars), respectively, for a photocopying machine satisfplease show all work and steps thank you :)
Answer #1

llt) = + 2 and Raja R (t) = stre a) for useful inze cct) = = R (t) +7/12 6-73 = st² et3 - e 9. = ln (+%) = -t ► t - Sen (50